### 12 Mar 2026 v1

# Random divergence-free drifts and the Onsager-Richardson threshold

Abstract.
We prove the absence of anomalous dissipation for passive scalars driven by random divergence-free autonomous vector fields in Hölder classes $C^{\alpha}(\mathbb{T}^{d})$ with $\alpha>\frac{1}{3}$, without any regularity assumptions on the passive scalar except that the initial data is fixed and bounded. The random field has to satisfy very mild conditions in two dimensions, while in higher dimensions it is assumed to have a special geometric structure. The proof relies on dimension-theoretic arguments, as opposed to commutator estimates. A consequence of these results is that anomalous regularization does not occur for this class of random vector fields.
Key words and phrases:
Anomalous dissipation, transport equation, Onsager’s conjecture, passive scalar transport, Sard property, renormalized solution
2020 Mathematics Subject Classification:
35Q49 (primary), 28A75, 35B33, 35D30, 35Q35, 47B80, 60H25, 76F25 (secondary)
1. Introduction
The phenomenon of anomalous dissipation, which is the nonvanishing of the viscous energy dissipation in the vanishing viscosity limit, occupies a central position in the mathematical theory of turbulence. It is therefore sometimes referred to as the zeroth law of turbulence. In a seminal 1949 note, Onsager [36] observed that weak solutions of the incompressible Euler equations conserve kinetic energy above Hölder regularity $1/3$, and may fail to conserve kinetic energy below Hölder regularity $1/3$. This prediction, now known as Onsager’s conjecture, has since been resolved: energy conservation holds for $C^{\alpha}$ solutions with $\alpha>1/3$ [30, 17, 28, 15], while non-conservative solutions exist for $\alpha<1/3$ [21, 22, 23, 12, 32, 13]. The exponent $1/3$ thus marks a sharp threshold separating rigidity from flexibility in nonlinear incompressible fluid dynamics. The problem of providing a rigorous example of anomalous dissipation in the vanishing viscosity limit for the (unforced) Navier–Stokes equations in 3D remains open.
A closely related question arises for passive scalar transport. Consider a scalar $\theta:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}$ which solves the advection–diffusion equation
$\partial_{t}\theta_{\varepsilon}+(v\cdot\nabla)\theta_{\varepsilon}=\varepsilon\Delta\theta_{\varepsilon},$ (1.1)
driven by a divergence-free velocity field $v:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ ($\mathbb{T}^{d}$ denotes the $d$-dimensional flat torus; note that in this paper we will consider autonomous vector fields). It is expected by physicists and confirmed by numerical studies that in a turbulent vector field the scalar variance dissipates at a non-vanishing rate as $\varepsilon\to 0$ [37, 26, 38, 39]. Moreover, in the physics literature, beginning with the works of Obukhov, Corrsin, and Yaglom [35, 18, 40] in 1949–1951, dimensional arguments analogous to the Kolmogorov’s 1941 theory of turbulence predict nontrivial flux relations for passive scalars in turbulent regimes. In particular, heuristic cascade theories suggest a relation of the form
$\alpha+2\beta=1,$
linking the Hölder regularity $C^{\alpha}(\mathbb{T}^{d})$ of the velocity field to an effective regularity $C^{\beta}(\mathbb{T}^{d})$ of the scalar. Such predictions are sometimes referred to as manifestations of anomalous regularization, whereby turbulent mixing produces effective smoothing properties that persist in the vanishing diffusivity limit. We will discuss this phenomenon in more detail in Section 4.3.
In mathematics, the existing results only scrape the surface of our understanding of anomalous dissipation or anomalous regularization for passive scalar transport. We are far from providing a rigorous foundation for a phenomenological theory. Only a few deterministic examples of particular vector fields that give rise to anomalous dissipation and occasionally to some version of the Obukhov-Corrsin-Yaglom balance have been built, in [27, 11, 16, 33, 10, 6, 29, 14] (see also [31]), and both anomalous dissipation and anomalous regularization have been recently established in the groundbreaking work [5] for rather general classes of random vector fields with a negative Hölder exponent, close to 1-1. The two effects are partially related as the occurrence of anomalous dissipation implies $\alpha+2\beta\leq 1$, cf. [27, 1]. Some of the aforementioned examples pertain to autonomous vector fields. We do not mention the results in stochastic PDEs pertaining to the Kraichnan model and its relatives as the requirements on roughness in time (white noise) position them too far from the context of the present note.
The present work establishes that in the autonomous setting $\alpha\leq 1/3$ is sharp, in the sense that random Hölder roughness above Onsager’s exponent $1/3$ almost surely prohibits anomalous scalar dissipation or anomalous regularization. We emphasise that we do not make Hölder regularity assumptions on the passive scalar $\theta$, our results can therefore be consider Onsager-supercritical. We consider probability measures on divergence-free autonomous average-free vector fields on the two-dimensional torus satisfying:
  1. (1)
    Hölder regularity $C^{\alpha}(\mathbb{T}^{2})$ with $\alpha>1/3$;
  2. (2)
    a mild nondegeneracy condition ensuring bounded density of the pointwise distribution.
Our main theorem shows that, under these assumptions, the associated transport equation possesses the DiPerna–Lions renormalization property almost surely. The latter property was introduced in the seminal work of DiPerna and Lions [25], where the authors proved it for any vector field with Sobolev regularity $W^{1,1}$. It is a consequence of their theory that then bounded weak solutions of the transport equation are “renormalized” (cf. Definition 2.1), uniqueness holds for the Cauchy problem, and anomalous dissipation is excluded. In fact, for any bounded initial data, solutions of (1.1) converge strongly, as diffusivity vanishes, to the unique inviscid transport solution, and thus scalar dissipation vanishes in the limit. An important extension of the renormalization property to $BV$ vector fields was later achieved by Ambrosio in [3], while the novel Lagrangian approach in [19] has spurred further extensions to other critical regularities, cf. [9]. The sharpness, at the deterministic level, of the thresold $\alpha=1$ in the Sobolev case is known since [24].
Let us state the details.
Definition 1.1.
Consider a bounded divergence-free (autonomous) vector field $v:\mathbb{T}^{d}\rightarrow\mathbb{R}^{d}$. We say that $v$ allows a dissipation anomaly if there is a bounded initial data $\theta_{in}\in L^{\infty}(\mathbb{T}^{d})$ with the following property. If $\theta_{\varepsilon}$ denotes, for any $\varepsilon>0$, the unique bounded solution of the advection-diffusion equation
$\left\{\begin{array}[]{ll}\partial_{t}\theta_{\varepsilon}+(v\cdot\nabla)\theta_{\varepsilon}=\varepsilon\Delta\theta_{\varepsilon}\\ \\ \theta_{\varepsilon}(\cdot,0)=\theta_{in}\end{array}\right$.\,, (1.5)
then the sequence $\{\theta_{\varepsilon}\}$ satisfies
$\limsup_{\varepsilon\downarrow 0}\,\varepsilon\int_{0}^{T}\int|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt>0$
for some positive $T$.
One main result of this note is the following.
Theorem 1.2 (Two-dimensional case).
Consider a probability measure $\mathbb{P}$ on the space of continuous divergence-free and average-free autonomous vector fields $v$ on $\mathbb{T}^{2}$. Assume that:
  • (a)
    There is $\alpha>\frac{1}{3}$ such that $v\in C^{\alpha}(\mathbb{T}^{2})$ $\mathbb{P}$-almost surely;
  • (b)
    There is a constant $C>0$ such that
    $\mathbb{P}(\{v:|v(x)-v_{0}|\leq r\})\leq Cr^{2}\qquad\forall r>0,\quad\forall x\in\mathbb{T}^{2}\,,\quad\mbox{and}\quad\forall v_{0}\in\mathbb{R}^{2}\,.$
Then $\mathbb{P}$-almost surely $v$ satisfies the DiPerna-Lions renormalization property (cf. Definition 2.1) and thus does not allow dissipation anomalies.
Note that the condition (b) is extremely mild, cf. Section 4.1, and in fact it can be further relaxed. The assumption that the vector field is average-free almost surely can be also easily removed, cf. Section 4.2: in fact allowing for the addition of a random constant drift only makes the result less interesting, cf. Remark 4.1. Our techniques allow for some extension to higher dimensions, although not quite as clean. We will formulate a general dimensional version in Section 2 below. Here we will limit ourselves to the case of three dimensions because it is easier to state, physically more relevant, and moreover relates the random vector field to the classical “Clebsch variables” of the fluid-dynamics literature.
Theorem 1.3 (Three-dimensional case).
Consider a probability measure $\mathbb{P}$ on the space of continuous $C^{1}$ maps $\phi=(\phi_{1},\phi_{2}):\mathbb{T}^{3}\to\mathbb{R}^{2}$ such that:
  • (a)
    There is $\alpha>\frac{1}{3}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{3};\mathbb{R}^{2})$ $\mathbb{P}$-almost surely;
  • (b)
    There is a constant $C>0$ such that
    $\mathbb{P}(\{\phi:|D\phi(x)-M_{0}|\leq r\})\leq Cr^{6}\qquad\forall r>0,\quad\forall x\in\mathbb{T}^{3}\,,\quad\mbox{and}\quad\forall M_{0}\in\mathbb{R}^{2\times 3}\,.$
Then $\mathbb{P}$-almost surely the vector field $v=\nabla\phi_{1}\times\nabla\phi_{2}$ satisfies the DiPerna-Lions renormalization property and thus does not allow dissipation anomalies.
The DiPerna-Lions renormalization property implies in fact much more, namely:
  • The corresponding transport equation $\partial_{t}\theta+(v\cdot\nabla)\theta=0$ has a unique bounded weak solution $\theta$ for any bounded initial data;
  • $\theta$ is constant on the trajectories of the unique regular Lagrangian flow of $v$ and its $L^{p}$ norms are constant in time;
  • The solutions $\theta_{\varepsilon}$ converge strongly in $C([0,\infty),L^{2}(\mathbb{T}^{d}))$ to $\theta$.
It is from these three properties that one immediately infers
$\lim_{\varepsilon\to 0}\varepsilon\int_{0}^{T}\int|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt=0\,$
for all $T>0$. This consequence of the third property mentioned above is very well-known (in Section 4.4 we give a quick derivation of it from precise statements which can be found in the many lecture notes on the DiPerna-Lions theory available in the literature, e.g. [4, 20]).
The proofs of Theorems 1.2 and 1.3 rely on a probabilistic Morse–Sard mechanism and a striking result of Alberti–Bianchini–Crippa [2] that shows that the renormalization property follows from a geometric condition: if the associated stream function (also referred to as the Hamiltonian) satisfies a Morse–Sard property (its set of critical values has zero Lebesgue measure), then bounded weak solutions are unique and renormalized. The nondegeneracy assumption on the underlying probability law implies, almost surely, a quantitative upper bound on the Hausdorff dimension of the critical set of the stream function. When $\alpha>1/3$, this dimension bound yields a Morse–Sard property via a deterministic covering argument. The theorem of Alberti–Bianchini–Crippa then implies the renormalization property of the vector field.
The exponent $1/3$ arises here from the inequality (cf. Lemma 3.2)
$1+\alpha>2-2\alpha,$
which governs the competition between Hölder regularity and Hausdorff dimension of the critical set. Thus the same threshold that appears in Onsager’s theory for the Euler equations reemerges in a purely geometric mechanism for passive scalar transport. The fact that the threshold exponent is identical is somewhat surprising. First, unlike the nonlinear energy-flux arguments from [17, 30, 28, 15] underlying Onsager’s theorem, the present mechanism is dimension-theoretic. Secondly, arguments supporting the anomalous dissipation for passive scalar transport and the Obukhov-Corrsin-Yaglom law of anomalous regularization, both in physics and in mathematics, rely on the dimensional analysis, energy cascade, and (sometimes) certain statistical homogeneities. To the best of our knowledge, none of them suggests that the exponent $\alpha=1/3$ might be critical in this scenario. In fact, one could argue that the existing deterministic examples, e.g., [27, 16, 33] extending to $\alpha<1$, suggest otherwise. The only exception is [6], which indeed restricts to $\alpha<1/3$, but their example is deterministic as well, and it is hard to say whether it is a hard restriction or an artefact of the method. One should also reiterate that our vector fields are autonomous (but so are the ones in [5] and [33]) and the results are probabilistically a.s., contrary to most of the positive results, with the notable exception of [5], where the authors are close to $\alpha=-1$. The latter work is moreover motivated by preexisting papers in the physics literature which use renormalization group ideas to analyze superdiffusivity, cf. for instance [8] (and also [7, 34]): neither in this literature there seems to be an indication that the exponent $\frac{1}{3}$ (see for instance the Richardson law in [8, p. 216]) is critical. We find it remarkable that the threshold for our arguments in Theorems 1.2 and 1.3 coincides with the Hölder regularity expected for a typical flow in homogeneous turbulence. We do not presently know if this is just a coincidence.
1.1. Acknowledgments
The research of the first author has been supported by the Cambridge Trust and the Cantab Capital Institute for Mathematics of Information. The research of the second and third authors has been supported by the Simons Foundation through the Simons Initiative on the Geometry of Flows (Grant Award ID BD-Targeted-00017375 CDeL, SM). The third author is also partially supported by the Simons Collaboration on Localisation of Waves, 563916 SM.
2. General theorem and reduction to the Morse-Sard property
Both Theorem 1.2 and Theorem 1.3 are special cases of the following more general one. In order to state it we first recall the Hodge $\star$ operator on (the flat torus) $\mathbb{T}^{d}$, which is a canonical way of mapping $k$-forms into (dk)(d-k)-forms. For our purposes we are interested only in the case $k=1$. If $x_{1},\ldots,x_{d}$ are the canonical coordinates on $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ the Hodge $\star$ maps the 1-form $\sum_{i=1}^{d}w_{i}dx_{i}$ to the d1d-1 form
$\star\sum_{i=1}^{d}w_{i}dx_{i}=\sum_{i=1}^{d}(-1)^{i}w_{i}dx_{1}\wedge\ldots\wedge dx_{i-1}\wedge\widehat{dx_{i}}\wedge dx_{i+1}\wedge\ldots\wedge dx_{d}\,.$
In particular the vector field $w=(w_{1},\ldots,w_{d})$ is divergence-free if and only if $d\star\sum_{i=1}^{d}w_{i}dx_{i}=0$.
The next concept which we introduce is the “renormalization property” of DiPerna-Lions.
Definition 2.1.
Let $v$ be a bounded divergence-free vector field on $\mathbb{T}^{d}$ (not necessarily autonomous). A bounded weak solution $\theta$ of the transport equation
$\left\{\begin{array}[]{ll}\partial_{t}\theta+(v\cdot\nabla)\theta=0,\\ \\ \theta(\cdot,0)=\theta_{in},\end{array}\right$. (2.1)
is a bounded function $\theta$ which satisfies
$\int_{0}^{\infty}\int_{\mathbb{T}^{d}}\theta(\partial_{t}\varphi+v\cdot\nabla\varphi)\,dx\,dt=-\int_{\mathbb{T}^{d}}\theta_{in}(x)\varphi(x,0)\,dx$ (2.2)
for every test function $\varphi\in C^{1}_{c}(\mathbb{T}^{d}\times\mathbb{R})$.
The vector field $v$ is said to have the DiPerna-Lions renormalization property if, for every bounded initial data $\theta_{in}\in L^{\infty}(\mathbb{T}^{d})$, every bounded weak solution $\theta$ of (2.1), and every test function $\beta\in C^{1}(\mathbb{R})$, the scalar $\beta(\theta)$ is a weak solution of (2.1) in the sense of equation (2.2) with initial data $\beta(\theta_{in})$.
Our main theorem is then the following.
Theorem 2.2.
Consider a probability measure $\mathbb{P}$ on the space of continuous, zero average, functions $\phi=(\phi_{1},\ldots,\phi_{d-1})\in C(\mathbb{T}^{d},\mathbb{R}^{d-1})$ satisfying the following two assumptions:
  • (a)
    There is $\alpha>\frac{1}{3}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ $\mathbb{P}$-almost surely;
  • (b)
    There is a constant $C$ such that
    $\mathbb{P}(\{\phi:|D\phi(x)-M_{0}|\leq r\})\leq Cr^{d(d-1)}\quad\forall r>0\,,\quad\forall x\in\mathbb{T}^{d}\,,\quad\mbox{and}\quad\forall M_{0}\in\mathbb{R}^{(d-1)\times d}\,.$ (2.3)
Then $\mathbb{P}$-almost surely, the divergence-free vector field $w$ such that
$\star\sum_{i}w_{i}dx_{i}=d\phi_{1}\wedge\ldots\wedge d\phi_{d-1}$ (2.4)
has the DiPerna-Lions renormalization property.
It is now easy to see that Theorem 1.2 and Theorem 1.3 are direct consequences of Theorem 2.2. In fact for $d=2$ the relation (2.4) is equivalent to $w=\nabla^{\perp}\phi_{1}$. Since $w$ is assumed, in Theorem 1.2, to be average-free, there is for each such $w$ a unique stream function $\phi_{1}:\mathbb{T}^{2}\to\mathbb{R}$ with zero average such that $w=\nabla^{\perp}\phi_{1}$: a probability measure as in Theorem 1.2 immediately translates into a probability measure as in Theorem 2.2. In the case of Theorem 1.3 a simple computation shows that (2.4) is in fact equivalent to $w=\nabla\phi_{1}\times\nabla\phi_{2}$. So far we have shown that the assumptions in Theorems 1.2 and 1.3 are in fact equivalent to those of Theorem 2.2 when $d=2$ and $3$, respectively.
The key to the proof of Theorem 2.2 is to show that $\phi$ satisfies the Morse-Sard property $\mathbb{P}$-almost surely and then use the deep result of Alberti, Bianchini, and Crippa in [2]. In particular, in Section 3 we will show the following.
Proposition 2.3.
Let $\mathbb{P}$ be a probability measure satisfying the same assumptions as in Theorem 2.2. Then $\mathbb{P}$-almost surely $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ has the Morse-Sard property, namely the set of its critical values $\{\phi(x):{\rm rank}\,(D\phi(x))\leq d-2\}$ is a (Lebesgue) null-set of $\mathbb{R}^{d-1}$.
With Proposition 2.3 at hand, we can use the results of [2] to conclude the renormalization property for any such $\phi$ which satisfies the Morse-Sard property. In the case of two-dimensional vector fields the statement follows immediately from Theorem 5.2 and the remark in Section 6.3 of [2]. In the case of higher-dimensional vector fields the statement follows from Section 6.2 and Section 6.3 in [2]. Note indeed that the components $\phi_{k}$ of the vector map $\phi$ satisfy $\nabla\phi_{k}\cdot w\equiv 0$, namely the condition required in [2, Section 6.2]. This orthogonality relation is easy to verify from the properties of the Hodge $\star$ operator. In fact for every vector field $z$ we have the relation
$(z\cdot w)dx_{1}\wedge\ldots\wedge dx_{d}=d\phi_{1}\wedge\ldots\wedge d\phi_{d-1}\wedge\sum_{i}z_{i}dx_{i}\,.$
In particular we conclude that, for any $k=1,\ldots,d-1$,
$(\nabla\phi_{k}\cdot w)dx_{1}\wedge\ldots\wedge dx_{d}=d\phi_{1}\wedge\ldots\wedge d\phi_{d-1}\wedge d\phi_{k}=0\,.$
3. Proof of the Morse-Sard property
The proof of Proposition 2.3 will be split into two steps. In the first step we will take advantage of probabilistic arguments. The statement is as follows.
Lemma 3.1.
Consider a probability measure $\mathbb{P}$ as in Proposition 2.3, and a function $\phi$ satisfying conditions (a) and (b) from Theorem 2.2. Then the Hausdorff dimension of the critical set
$Z:=\{{\rm rank}\,(D\phi)\leq d-2\}$ (3.1)
is, $\mathbb{P}$-almost surely, at most $d-2\alpha$.
We then use a purely deterministic argument in the second part of the proof.
Lemma 3.2.
Let $\phi\in C^{1,\alpha}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ be a function such that the Hausdorff dimension of its critical set $Z$ is at most $d-2\alpha$. If $\alpha>\frac{1}{3}$, then the set $\phi(Z)$ is a (Lebesgue) null-set of $\mathbb{R}^{d-1}$.
Clearly Proposition 2.3 is a direct consequence of Lemmas 3.1 and 3.2.
Proof of Lemma 3.1.
In all the arguments, we always consider functions $\phi$ in the regularity class $C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$.
Step 1. We consider the Euclidean space $\mathbb{R}^{(d-1)\times d}$ of $(d-1)\times d$ real matrices and denote by $W$ the closed subset of matrices of rank at most d2d-2. If we denote by ${\rm dim}_{H}(E)$ the Hausdorff dimension of a set $E$ and introduce the set
$\mathcal{Y}:=\{\phi:{\rm dim}_{H}((D\phi)^{-1}(W))>d-2\alpha\}\,$,
then we can rephrase the claim of the theorem as $\mathbb{P}(\mathcal{Y})=0$. On the other hand $\mathcal{Y}$ is the countable union of
$\mathcal{Y}_{N}:=\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad\mathcal{H}^{d-2\alpha+1/N}((D\phi)^{-1}(W\cap B_{N}))>0\}\,$, (3.2)
where $B_{N}\subset\mathbb{R}^{(d-1)\times d}$ denotes the closed ball of radius $N\in\mathbb{N}$ (centered at the origin) and $\mathcal{H}^{\beta}$ is the $\beta$-dimensional Hausdorff measure. In particular it suffices to show that
$\mathbb{P}(\mathcal{Y}_{N})=0\qquad\forall N\in\mathbb{N}\setminus\{0\}\,.$ (3.3)
Step 2. From now on we fix a positive natural number $N$ and aim at proving (3.3). First of all we recall that $W$ is a real-algebraic subset of $\mathbb{R}^{(d-1)\times d}$ of codimension $2$. Its intersection with the closed ball $B_{N}$ has, therefore, finite Minkowski (d(d1)2)(d(d-1)-2)-dimensional content. In other words there is a constant $C$ (depending on both $W$ and $N$ but not on $\delta$) such that, for every $\delta>0$,
$|\{A\in\mathbb{R}^{(d-1)\times d}:{\rm dist}\,(A,W\cap B_{N})\leq\delta\}|\leq C\delta^{2}\,.$
Consider now the cubical decomposition of $\mathbb{R}^{d(d-1)}$ in cubes $Q:=\delta M+[0,\delta)^{d(d-1)}$, for $M$ ranging in the lattice $\mathbb{Z}^{d(d-1)}$. The cubes are chosen so to be pairwise disjoint. We then let $\mathcal{Q}_{\delta}$ be the subcollection of such cubes which intersect the set $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta\}$. The cardinality of this collection is controlled by $C\delta^{-d(d-1)+2}$, where $C$ is another constant independent of $\delta$: this follows because the cubes are pairwise disjoint and their union is contained in $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta(1+\sqrt{d})\}$.
Observe next that, by assumption (b) of Theorem 2.2, for any $\delta>0$ and any fixed $Q$ in $\mathcal{Q}_{\delta}$ (for $x\in\mathbb{T}^{d}$)
$\mathbb{P}(\{\phi:D\phi(x)\in Q\})\leq C\delta^{d(d-1)}\,.$
In particular, summing over all the cubes $Q$ in $\mathcal{Q}_{\delta}$, we conclude immediately the bound
$\mathbb{P}(\{\phi:{\rm dist}\,(D\phi(x),W\cap B_{N})\leq\delta\})\leq C\delta^{2}\,.$ (3.4)
Step 3. We next fix a natural number $k$, we subdivide $\mathbb{T}^{d}$ in $2^{-kd}$ closed nonoverlapping cubes $Q$ of sidelength $2^{-k}$, and for each $\phi\in C^{1,\alpha}$ we count the number $n(k,\phi)$ of cubes in the subdivision which intersect $(D\phi)^{-1}(W\cap B_{N})$. Let us enumerate the cubes of the subdivision as $Q_{i}$ and note that
$n(k,\phi)=\sum\mathbf{1}_{E_{i}}(\phi)$ (3.5)
where $E_{i}$ is the set of $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ for which there is a point $y\in Q_{i}$ such that $D\phi(y)\in W\cap B_{N}$.
Note that, if $x_{i}$ is the center of the cube $Q_{i}$ and $\|D\phi\|_{C^{1,\alpha}}\leq N$, then
$|D\phi(y)-D\phi(x_{i})|\leq N|x-y_{i}|^{\alpha}\leq N(\sqrt{d}\,2^{-k-1})^{\alpha}\,.$ (3.6)
Thus, if we introduce the set
$F_{i}:=\{\phi:{\rm dist}\,(D\phi(x_{i}),W\cap B_{N})\leq N(\sqrt{d}\,2^{-k-1})^{\alpha}\}\,$,
we infer from estimate (3.6) that $E_{i}\cap\{|D\phi\|_{C^{1,\alpha}}\leq N\}\subset F_{i}$ and
$n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}}(\phi)\leq\sum_{i}\mathbf{1}_{F_{i}}(\phi)\,.$ (3.7)
By inequality (3.4), by choosing $\delta=N(\sqrt{d}\,2^{-k-1})^{\alpha}$ we have $\mathbb{P}(F_{i})\leq CN^{2}2^{-2k\alpha}$ and therefore, integrating (3.7) and by taking into account that there are $2^{kd}$ distinct cubes in the cubical decomposition of $\mathbb{T}^{d}$ considered above, we conclude immediately
$\int_{\|D\phi\|_{C^{1,\alpha}}\leq N}n(k,\phi)\,d\mathbb{P}(\phi)\leq\sum_{i}\mathbb{P}(F_{i})\leq CN^{2}2^{(d-2\alpha)k}\,.$ (3.8)
We next use Chebyshev’s inequality for the function $n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}}$ to estimate the probability of the event
$\mathcal{Y}_{N,k}:=\left\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad n(k,\phi)\geq 2^{(d-2\alpha+1/(2N))k}\right\}\,$
and get therefore
$\mathbb{P}(\mathcal{Y}_{N,k})\leq 2^{-(d-2\alpha+1/(2N))k}\mathbb{E}\big[n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}(\phi)\big]}\leq CN^{2}2^{-k/(2N)}\,.$ (3.9)
By the Borel-Cantelli lemma, the set
$\mathcal{Y}^{\prime}_{N}:=\bigcap_{j=1}^{\infty}\bigcup_{k\geq j}\mathcal{Y}_{N,k}$
is a set of measure zero.
We now claim that $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, which would conclude the proof. In fact, assume $\phi\in\mathcal{Y}_{N}$. If $\phi$ is in the complement of $\mathcal{Y}^{\prime}_{N}$ then necessarily there would be a sequence of natural numbers $k(i)\to\infty$ with the property that $\phi\not\in\mathcal{Y}_{N,k(i)}$ for every $i$. But, given the definition of $\mathcal{Y}$, $\|\phi\|_{C^{1,\alpha}}\leq N$. This means that, for every $i$, $n(k(i),\phi)<2^{(d-2\alpha+1/(2N))k(i)}$. Therefore we can cover $(D\phi)^{-1}(W\cap B_{N})$ with less then $2^{(d-2\alpha+1/(2N))k(i)}$ cubes of sidelength $2^{-k(i)}$. We then conclude, from the definition of the $\mathcal{H}^{\beta}_{\infty}$ premeasure, that
$\mathcal{H}^{d-2\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))\leq\omega_{d-2\alpha+N^{-1}}2^{-k(i)/(2N)}\left(\frac{\sqrt{d}}{2}\right)^{d-2\alpha+N^{-1}}\,.$ (3.10)
But letting $i\uparrow\infty$ we would then conclude $\mathcal{H}^{d-2\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))=0$. Since $\mathcal{H}^{\beta}_{\infty}$-null sets and $\mathcal{H}^{\beta}$-null sets coincide, we then would be in contradiction with the definition of $\mathcal{Y}_{N}$, which requires $\mathcal{H}^{d-2\alpha+N^{-1}}((D\phi)^{-1}(W\cap B_{N}))>0$. Therefore we have $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, from which it follows that both $\mathbb{P}(\mathcal{Y}_{N})$ and $\mathbb{P}(\mathcal{Y})$ are equal to zero. ∎
Proof of Lemma 3.2.
Since $Z$ has Hausdorff dimension at most $d-2\alpha$ (as was proved in Lemma 3.1), for every positive $\delta$ we have $\mathcal{H}^{d-2\alpha+\delta}(Z)=0$. We fix here a $\delta$ such that
$3\alpha\geq 1+\delta\,.$ (3.11)
From the definition of Hausdorff measure it follows that for every positive $\varepsilon\leq 1$ there is a cover of $Z$ with sets $U_{k}$ such that
$\sum_{k}({\rm diam}\,(U_{k}))^{d-2\alpha+\delta}<\varepsilon\,.$ (3.12)
We then discard any $U_{k}$ which does not intersect $Z$ and for the remaining ones we pick a point $x_{i}\in U_{k}\cap Z$. Set $r_{i}:={\rm diam}\,(U_{i})$ and conclude that the closed balls $B_{r_{i}}(x_{i})$ cover $Z$ and satisfy (by rewriting equation (3.12))
$\sum_{k}r_{k}^{d-2\alpha+\delta}<\varepsilon\,.$ (3.13)
Next we recall that $\phi$ is Lipschitz and thus $\phi(U_{i})\subset B_{C_{0}r_{i}}(\phi(x_{i}))$, where $C_{0}={\rm Lip}\,(\phi)$. Moreover, the rank of the linear map $L_{i}:=D\phi(x_{i})$ is at most d2d-2, as $x_{i}\in Z$: this allows us to choose a (d2)(d-2)-dimensional vector space $V_{i}$ with the property that ${\rm Im}\,(L_{i})\subset V_{i}$.
Select then a unit normal vector $\nu_{i}$ which is orthogonal to $V_{i}$ and observe that, by Lagrange’s mean-value theorem, for every $y\in B_{r_{i}}(x_{i})$ there is a point $\xi$ on the segment $x_{i},y]$ such that
$\nu_{i}\cdot(\phi(y)-\phi(x_{i}))=\nu_{i}\cdot(D\phi(\xi)\cdot(y-x_{i}))\,.$
Since $\nu_{i}\cdot(D\phi(x_{i})\cdot(y-x_{i}))=0$ (due to the orthogonality of $\nu_{i}$ with respect to $V_{i}$), we then conclude that
$|\nu_{i}\cdot(\phi(y)-\phi(x_{i}))|$ $=|\nu_{i}\cdot((D\phi(\xi)-D\phi(x_{i}))\cdot(y-x_{i}))|\leq[D\phi]_{0,\alpha}|\xi-x_{i}|^{\alpha}|y-x_{i}|$
$\leq[D\phi]_{0,\alpha}r_{i}^{1+\alpha}\,.$
In particular the orthogonal projection of $\phi(B_{r_{i}}(x_{i}))$ onto the orthogonal complement $V_{i}^{\perp}$ of $V_{i}$ is contained in a closed segment $\sigma_{i}$ of length $2[D\phi]_{0,\alpha}r_{i}^{1+\alpha}$ centered on the projection of $\phi(x_{i})$ onto $V_{i}^{\perp}$. If we denote by $z_{i}$ the projection of $\phi(x_{i})$ onto $V_{i}$, the arguments above prove the inclusion
$\phi(B_{r_{i}}(x_{i}))\subset(B_{C_{0}r_{i}}(z_{i})\cap V)+\sigma_{i}=:W_{i}\,.$
On the other hand, since $B_{C_{0}r_{i}}(z_{i})\cap V$ is a closed (d2)(d-2)-dimensional disk, the set $W_{i}$ can be covered by $N_{i}\leq\bar{C}r_{i}^{-\alpha(d-2)}$ balls (of $\mathbb{R}^{d-1}$) with radius $r_{i}^{1+\alpha}$, where $\bar{C}$ depends only on $C_{0}$, $D\phi]_{0,\alpha}$, and $d$. The union over $i$ of all these balls covers therefore $\phi(Z)$ and we can write
$|\phi(Z)|$ $\leq\omega_{d-1}\sum_{i}N_{i}(r_{i}^{1+\alpha})^{d-1}\leq\bar{C}\omega_{d-1}\sum_{i}r_{i}^{(1+\alpha)(d-1)-\alpha(d-2)}$
$=\bar{C}\omega_{d-1}\sum_{i}r_{i}^{d-1+\alpha}\,.$ (3.14)
Observe that $d-1+\alpha\geq d-2\alpha+\delta$ because of nequality (3.11). Moreover, since $\varepsilon\leq 1$, inequality (3.13) implies that $r_{i}\leq 1$ for all $i$’s and in particular $r_{i}^{d-1+\alpha}\leq r_{i}^{d-2\alpha+\delta}$. Thus (3.13) and (3.14) immediately imply
$|\phi(Z)|\leq\bar{C}\omega_{d-1}\varepsilon\,.$
Since $\varepsilon$ can be chosen arbitrarily small, we then conclude that $\phi(Z)$ has zero Lebesgue measure, which is what we had to show. ∎
4. Remarks
In this section we take Theorem 1.2 as base of our discussions. Analogous statements hold for Theorem 1.3 and Theorem 2.2, but we leave the details to the reader.
4.1. Random vector fields
First of all, we observe that assumption (b) in Theorem 1.2 is equivalent to the probability distribution of the ($\mathbb{R}^{2}$-valued) random variable $v(x)$ having bounded density. This is an extremely mild assumption, which is satisfied in a number of standard choices of $\mathbb{P}$. Consider for instance the usual construction of random vector fields as follows
[…]
$H=\left\{v\in L^{2}(\mathbb{T}^{2}):{\rm div}\,v=0\quad\mbox{and}\quad\int_{\mathbb{T}^{2}}v\,dx=0\right\}\,.$
[…]
$C(x)=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\sigma_{k}v_{k}(x)\otimes v_{k}(x)\,.$
Assumption (b) amounts therefore to having a uniform lower bound $C(x)\geq C_{0}{\rm Id}$ in the sense of positive definite matrices. If we take classical bases such as Fourier or (several bases of) wavelets, the uniform lower bound is implied by the positivity of the $\sigma_{k}$’s. In fact the positivity of them for a handful of low frequencies is already enough.
Furthermore, a simple inspection of the proof shows that indeed condition (b) is only needed for the special value $v_{0}=0$, while in the general case of Theorem 1.3 the condition (b) is only needed for $M_{0}$’s which have rank at most d2d-2.
4.2. Adding constant drifts
The assumption, in Theorem 1.2, that the average $\langle v\rangle$ of the random vector field $v$ vanishes almost surely can be easily removed. Here we quickly explain how. First of all, the arguments of Lemma 3.1 can be used verbatim to show that, $\mathbb{P}$-almost surely, the set $\{v=0\}$ has dimension at most $2-2\alpha$: the proof only uses the assumptions (a) and (b). Fix now such a $v$. The claim is that it satisfies the conclusions of the theorem because in fact it has the DiPerna-Lions renormalization property: the assumption $\langle v\rangle=0$ is not essential in this step either. First of all Lemma 3.2 can be used verbatim to show that, in every subset $\Omega$ where a function $\psi$ with $\nabla\psi^{\perp}=v$ exists, the critical values of $\psi$ has measure zero. Since this will hold for every $\Omega$ which is simply connected, the results in [2] can then be used to show that, when $\Omega$ is simply connected and $\theta$ and $\beta$ are as in Definition 2.1, then $\beta(\theta)$ will be a distributional solution of the transport equation (2.1) in $\Omega\times[0,\infty)$ with initial data $\beta(\theta_{in})$. But a simple partition of unity allows then to show that $\beta(\theta)$ is in fact a distributional solution on the whole torus.
Generalizing the above to Theorem 2.2 is rather straightforward: we just need to make $\mathbb{P}$ a probability distribution on the space of $C^{1}$ maps $\phi:\mathbb{R}^{d}\to\mathbb{R}^{d-1}$ which are the sums of a $1$-periodic function and a linear map. The differential of each component $\phi_{i}$ is then a periodic $1$-form and induce therefore a closed $1$-form $\omega_{i}$ on $\mathbb{T}^{d}$. In particular the conclusion of Theorem 2.2 will then hold for the divergence-free vector field $w$ (which needs not be mean-free) such that $\star\sum_{i}w_{i}dx_{i}=\omega_{1}\wedge\ldots\wedge\omega_{d-1}$.
Remark 4.1.
The addition of a sufficiently large constant to a $2$-dimensional $C^{1}$ divergence-free autonomous vector field makes it a nowhere zero vector field. It follows then immediately from our discussion that $w$ has the renormalization property: this of course a very simple case of the general theory developed in [2]. Consider now a probability measure $\mathbb{P}$ on divergence-free autonomous vector fields with the following two properties:
  • $v\in C^{0}$ $\mathbb{P}$-almost surely;
  • The average $\langle v\rangle$ of $v$ and its average-free part $v-\langle v\rangle$ are independent;
  • $\mathbb{P}(\{v:|\langle v\rangle|\geq M\})>0$ for every constant $M$.
Then there is a nonzero probability $\mathbb{P}$ that $v$ has no zeros, and thus a nonzero probability that it has the renormalization property. Note that the probability is, potentially, rather small but that no Hölder regularity of $v$ is needed. Note that the fact that in Theorems 1.2, 1.3 and 2.2 we restrict to zero average vector fields means that they have to be zero somewhere.
4.3. Yaglom’s law
A widely accepted theory in the turbulence literature is that there is, in some statistical sense, a relation between the regularity of the advecting vector field $v$ and the passive scalar $\theta_{\varepsilon}$ in (1.5). Roughly speaking a $C^{\alpha}$ regularity for $v$ should correspond to a uniform $C^{\beta}$ regularity of $\theta_{\varepsilon}$ for $\beta$ satisfying $\alpha+2\beta=1$. Some authors in the literature call this “anomalous regularization”. More generally the “law” is known as Yaglom’s Law or the Obukhov-Corrsin relation. Clearly a deterministic statement of the latter cannot hold. However, Theorem 1.2 (and, to a lesser extent, also Theorem 2.2) show that even at a statistical level such a statement cannot hold when $v$ is autonomous and $\alpha>\frac{1}{3}$. In fact, because of the DiPerna-Lions renormalization property, when the conclusion of Theorem 2.2 applies and the initial data $\theta_{in}$ is bounded, the solutions $\theta_{\varepsilon}$ of (1.5) converge (strongly in $C([0,T],L^{2})$) to the corresponding unique solution $\theta$ of (2.1). Moreover $\theta$ is constant along the trajectories of the unique regular Lagrangian flow $\Phi$ of the vector field $v$. Therefore the distribution function of $a\mapsto|\{\theta>a\}|$ will remain exactly the same. This immediately obstructs any regularization mechanism if, for instance, the values taken by $\theta_{in}$ is a discrete set with more than one element: $\theta(\cdot,t)$ will be forced to remain discontinuous on a set of positive $\mathcal{H}^{1}$ measure.
Additionally, Remark 4.1 suggests that a proper framework to study random drifts is likely that of average zero vector fields.
4.4. From the renormalization property to the absence of dissipation anomalies
We provide for the reader’s convenience a quick derivation of the following proposition from statements which can be found in the literature on the transport equation.
Proposition 4.2.
Consider a bounded divergence-free vector field $v:\mathbb{T}^{d}\times[0,T]\to\mathbb{R}^{d}$ which satisfies the renormalization property . Let $\theta_{in}$ be a bounded initial datum and for any $\varepsilon>0$ we denote by $\theta_{\varepsilon}$ the unique bounded solution of equation (1.5). Then we have
$\lim_{\varepsilon\downarrow 0}\,\varepsilon\int_{0}^{T}\int_{\mathbb{T}^{d}}|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt=0$ (4.1)
and $\theta_{\varepsilon}$ converges strongly in $C([0,T],L^{2}(\mathbb{T}^{d}))$ to the unique bounded weak solution of (2.1).
Proof.
First of all the existence and uniqueness of the bounded distributional solution $\theta_{0}$ of (2.1) is the main result of the DiPerna-Lions theory and the reader can consult e.g. [20, Theorem 3.3 & Proposition 3.6]. The strong continuity of $t\mapsto\theta(\cdot,t)$ (up to a suitable choice of representative) is another well-known fact and the reader can consult e.g. [4, Remark 25]. By the maximum principle $\theta_{\varepsilon}$ is uniformly bounded in $L^{\infty}((0,T);L^{\infty}(\mathbb{T}^{d}))$ and so, up to subsequence, it converges weakly-* to some bounded function $\theta$. We immediately see (passing to the limit in the weak formulation (2.2)) that $\theta$ is a bounded distributional solution of equation (2.1) and the aforementioned uniqueness implies that $\theta=\theta_{0}$. Fix a smooth test function $\varphi\in C^{\infty}(\mathbb{T}^{d})$ and define $\Phi_{\varepsilon}(t):=\int_{\mathbb{T}^{d}}\varphi(x)\theta_{\varepsilon}(x,t)dx$. Multiplying equation (1.5) with the function $\varphi(x)$ and integrating in space, we immediately obtain that
$\frac{d}{dt}\Phi_{\varepsilon}(t)=\int_{\mathbb{T}^{d}}\theta_{\varepsilon}(x,t)\left[v(x,t)\cdot\nabla\varphi(x)+\varepsilon\Delta\varphi(x)\right]\,dx\,.$
In particular the functions $t\mapsto\Phi_{\varepsilon}(t)$ are equi-Lipschitz. By the Ascoli-Arzelà Theorem and the uniqueness (and time continuity) of the limit $\theta_{0}$, we conclude that
$\lim_{\varepsilon\downarrow 0}\Phi_{\varepsilon}(t)=\int_{\mathbb{T}^{d}}\theta_{0}(x,t)\varphi(x)\,dx\,.$
The arbitrariness of the test function $\varphi$ and the boundedness of $\theta_{\varepsilon}(\cdot,t)$ implies therefore that $\theta_{\varepsilon}(\cdot,t)\rightharpoonup\theta_{0}(\cdot,t)$ in $L^{2}(\mathbb{T}^{d})$ for every $t\in[0,T]$. In particular, recalling that
$\int_{\mathbb{T}^{d}}\theta^{2}_{\varepsilon}(x,T)\,dx=\int_{\mathbb{T}^{d}}\theta_{in}^{2}(x)\,dx-2\varepsilon\int_{0}^{T}\int_{\mathbb{T}^{d}}|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt$ (4.2)
and using the lower semicontinuity of the $L^{2}$ norm under weak convergence, we conclude
$\int_{\mathbb{T}^{d}}\theta_{0}^{2}(x,T)\,dx\leq\liminf_{\varepsilon\downarrow 0}\int_{\mathbb{T}^{d}}\theta_{\varepsilon}^{2}(x,T)\,dx\leq\int_{\mathbb{T}^{d}}\theta_{in}^{2}(x)\,dx\,.$ (4.3)
However, the renormalization property gives that $\theta_{0}^{2}$ is the unique bounded weak solution of (2.1) with initial data $\theta_{in}^{2}$ and testing the equation against functions of the type $\varphi(x,t)=\varphi(t)$ we conclude that $\|\theta_{0}(\cdot,t)\|_{L^{2}}^{2}$ is constant. So the inequalities in (4.3) are actually equalities and in fact $\theta_{\varepsilon}(\cdot,T)$ converges strongly in $L^{2}(\mathbb{T}^{d})$ to $\theta(\cdot,T)$. Coupling this information with (4.2) we conclude (4.1). Note also that the argument immediately implies that $\theta_{\varepsilon}(\cdot,t)$ converges strongly to $\theta_{0}(\cdot,t)$ in $L^{2}(\mathbb{T}^{d})$ for every $t\in[0,T]$. To upgrade pointwise convergence to uniform convergence we observe that we already know the uniform convergence in $L^{2}(\mathbb{T}^{d})$ endowed with the weak topology and that the functions $t\mapsto\|\theta_{\varepsilon}(\cdot,t)\|_{L^{2}}$ are converging to $t\mapsto\|\theta_{0}(\cdot,t)\|_{L^{2}}\equiv\|\theta_{in}\|_{L^{2}}$ uniformly. ∎